# FTFT: Efficient and Robust Fine-Tuning by Transferring Training Dynamics

### v2

@@ 6 Our Approach: FTFT
Building on our insights presented in the previous sections, we now propose a novel fine-tuning approach based on dataset cartography, Fine-Tuning by transFerring Training dynamics (FTFT). Compared with the original dataset cartography approach, FTFT integrates two crucial changes: (1) using more efficient reference models, which not only result in comparable DMs but also enhance efficiency, and (2) implementing aggressive early stopping during the main model training. Such aggressive early stopping is possible because models trained with data instances selected using DMs already achieve strong performance with substantially fewer training steps (§5).
@@
We show our results on NLI in Table 1 (FTFT) and on HSD in Table 5 (Appendix B). We use an early stopping patience (i.e. the number of checkpoints without improvement before stopping), based on the average dev set performance of the OOD datasets. We use a relatively small , because it can stop the training earlier, leading to higher efficiency. We also show the computational cost of each method (Compute), in which we have taken the training of both the reference model and the main model into account, including the extra checkpoints after the best performance is achieved.
@@
We make three observations. First, FTFT achieves consistent robustness improvement over ERM, indicated by its strong performance on most OOD datasets (note again that the hyper-parameters are optimized for ERM). We also include the results of ERM with early stopping (ERM(ES)) as a baseline to illustrate that our aggressive early stopping strategy is only effective when training with data instances selected by DMs: ERM(ES) achieves worse performance than ERM on all OOD datasets. Second, FTFT achieves substantial higher training efficiency than other methods. For example, on NLI, training using FTFT and as reference models only costs 51.57% of the training cost of ERM, which is equivalent to 25.78% of the training cost of the original dataset cartography approach. Third, our early stopping strategy is effectively aggressive. Specifically, FTFT with all three reference models in Table 1 achieves their best performance within just 1/3 of the optimal training steps of ERM.
7 Conclusions
@@
Fine-tuned PLMs have been shown to be vulnerable to OOD inputs. Although dataset cartography can improve model robustness (Swayamdipta et al., 2020; Sar-Shalom & Schwartz, 2023), it is computationally expensive. In this paper, we have presented FTFT, a novel approach for fine-tuning PLMs which yields both better efficiency and better robustness over ERM (§6). FTFT is built on dataset cartography, based on two observations: (1) reference model training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2), and (2) main models trained using data instances selected by reference model training dynamics achieve higher learning speed (i.e. substantially better performance with reduced number of training steps). We have also discussed the conditions for successful FTFT runs (§4.3). We believe that FTFT will be an important tool for future researchers and practitioners to perform efficient model fine-tuning, especially in situations where robustness is essential. We also believe our study of lightweight reference models will benefit future research on building efficient supervised data pruning metrics, which are shown to outperform their unsupervised counterparts (Sorscher et al., 2022).
@@
Our work opens up several directions for future research. First, we have observed that effective reference models identify more instances as easy. Further controlled experiments are needed to build a more rigorous protocol for choosing reference models. We believe the results presented here provide a good foundation for such further work. Second, we have empirically identified the transferability of training dynamics to selected training instances. Future studies are needed to build the theoretical foundations of both data cartography itself and the feasibility of such transfers. Third, we have only developed FTFT for classification tasks. Future studies can extend our work to other tasks, such as generation (e.g., question answering) and self-supervised tasks (e.g., language modeling).
@@ A.1 Experimental Setup
For training all models, we use AdamW (Loshchilov & Hutter, 2019) as the optimizer with a batch size of 32. We also use a linear learning rate scheduler with 10% warmup. For fine-tuning the small and base versions of both DeBERTaV3 and ELECTRA, as well as TinyBERT, RoBERTa, and BERT, we use a learning rate of 2e-5, following the suggestions from the original papers. We use a smaller learning rate of 1e-5 for and , because training larger models with lower learning rates are observed to be more stable (Mosbach et al., 2021; Du & Nguyen, 2023). However, a few failed training runs (i.e., training fails to converge and the resulting model performs worse than the majority class baseline Mosbach et al., 2021) are still observed during the training of . We excluded these runs from our results.

### v3

@@ 6 Our Approach: FTFT
Building on our insights presented in the previous sections, we propose a novel fine-tuning approach based on dataset cartography (Swayamdipta et al., 2020), Fine-Tuning by transFerring Training dynamics (FTFT). Compared with the original dataset cartography approach, FTFT integrates two crucial improvements: (1) using more efficient reference models, which not only result in comparable DMs but also enhance efficiency, and (2) implementing aggressive early stopping during the main model training. Such aggressive early stopping is possible because models trained with data instances selected using DMs already achieve strong performance with substantially fewer training steps (§5). In summary, FTFT consists of three steps: (1) train an efficient reference model on the full dataset, (2) compute the ambiguity of each instance based on the standard deviation of the correct class probability across epochs, and (3) select the top most ambiguous instances to retrain a larger main model, while applying aggressive early stopping.
@@
We show our results on NLI in Table 1 (FTFT) and on HSD in Table 5 (Appendix B). We use an early stopping patience (i.e. the number of checkpoints without improvement before stopping), based on the average dev set performance on the OOD datasets. We use a relatively small , as stopping the training earlier will lead to higher efficiency. We also show the computational cost of each method (Compute), in which we have taken the training of both the reference model and the main model into account, including the extra checkpoints after the best performance is achieved.
@@
We have three key observations. First, FTFT achieves consistent robustness improvements over ERM, indicated by its strong performance on most OOD datasets (note again that the hyper-parameters are optimized for ERM). We also include the results of ERM with early stopping (ERM(ES)) as a baseline to illustrate that our aggressive early stopping strategy is only effective when training with data instances selected by DMs: ERM(ES) achieves worse performance than ERM on all OOD datasets. Second, FTFT substantially improves the training efficiency of models. For example, on NLI, training using FTFT and as the reference model only costs 51.57% of ERM’s training cost, which is equivalent to 25.78% of the training cost of the original dataset cartography approach. Third, our early stopping strategy is effectively aggressive. Specifically, all FTFT results in Table 1 are achieved using less than 1/3 of the optimal training steps needed for ERM.
7 Conclusion
@@
Fine-tuned PLMs have shown to be vulnerable to OOD input. Although dataset cartography can improve model robustness (Swayamdipta et al., 2020; Sar-Shalom and Schwartz, 2023), it is computationally expensive. In this paper, we have presented FTFT, a novel approach for fine-tuning PLMs which yields both better efficiency and better robustness over ERM (§6). FTFT is built on dataset cartography, based on two observations: (1) reference model training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2), and (2) main models trained using data instances selected using DMs learn faster (i.e. they have substantially better performance with reduced number of training steps). We believe that FTFT will be an important tool for future researchers and practitioners to perform efficient model fine-tuning, especially in situations where robustness is essential.
8 Limitations
@@
We identify three limitations from this work. These limitations open up promising directions for future research. First, we have observed that effective reference models identify more instances as easy. More controlled experiments are needed to build a detailed protocol for choosing efficient but strong enough reference models. Our results provide a good foundation for such further work. Second, we have empirically demonstrated the transferability of training dynamics to select training instances. Future studies are needed to build the theoretical foundations of both data cartography itself and the feasibility of such transfers. Third, we have only developed FTFT for classification tasks. Future studies should extend FTFT to generation tasks, e.g., instruction following and language modeling.
Acknowledgments
We thank the anonymous reviewers for their helpful feedback. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grants no. EINF-10999 and EINF-5693.
@@ A.1 Experimental Setup
For training all models, we use AdamW (Loshchilov and Hutter, 2019) as the optimizer with a batch size of 32. We also use a linear learning rate scheduler with 10% warmup. For fine-tuning the small and base versions of both DeBERTaV3 and ELECTRA, as well as TinyBERT, RoBERTa, and BERT, we use a learning rate of 2e-5, following the suggestions from the original papers. We use a smaller learning rate of 1e-5 for and , because training larger models with lower learning rates are observed to be more stable (Mosbach et al., 2021; Du and Nguyen, 2023). However, a few failed training runs (i.e., runs where the training fails to converge and where the resulting model performs worse than the majority class baseline (Mosbach et al., 2021)) still occurred during the training of . We excluded these runs from our results.
@@ A.2 Comparison of Training Costs
We estimate the training FLOPs by their models sizes (i.e. the number of parameters), following Kaplan et al. (2020) (, where is the number of FLOPs, is the number of parameters, and is the dataset size). The results are in Table 3. We prefer theoretical estimation over practical measurement, because (1) the results are less influenced by noises, (2) it considers both forward and backward propagation, and (3) it has been shown effective and widely used (Kaplan et al., 2020; Hoffmann et al., 2022). In practice, we also measured the actual FLOPs usage in a forward run using Sovrasov (2018), and the results are very close to our estimation using model sizes.
@@ Appendix B Additional Results
| When Do Transfers Fail: Using Reference Models of Different Capabilities | ||||||||||